OpenStudy (anonymous):

Rewrite the equation below in standard form -3x+6y=12. Can you please also explain how to do it?

OpenStudy (anonymous):

These are the answer choices A. 3x-6y=12 B. 3x-6y=-12 C. -x+2y=4 D. x-2y=-4

OpenStudy (mathlegend):

In standard form the lead coefficient needs to be a positive. In this case it is negative. So multiply every term by -1.

OpenStudy (mathlegend):

-1(-3x+6y=12) What do yo uget @HUE-1?

OpenStudy (anonymous):

3x-6y=-12?

OpenStudy (mathlegend):

Great :)

OpenStudy (anonymous):

So the answer would be a positive 12 or negative?

OpenStudy (mathlegend):

It is a negative 12.